Gilmore, Inc., had equity of $145,000 at the beginning of the year. At the end of the year, the company had total assets of $What is the internal growth rate using beginning of period assets?
Answer #1

Beginning of period assets= End of period assets-(Net Income-Dividend)

= 210000-(27000-5800)

=188800

IGR = ROA*b

= (Net Income/Beginning assets)*(Addition to retained earnings/Net Income)

= (27000/188800)*((27000-5800)/27000)

=0.143008*0.785185

=11.23%